OpenStudy (anonymous):

multiplicative inverse of 1/ 3+i

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

the answer is \[3+i\]

OpenStudy (anonymous):

thanks...could you explain that for me?

OpenStudy (anonymous):

because \[\frac{1}{3+i}\times (3+i)=1\]

OpenStudy (saifoo.khan):

Surely he can!

OpenStudy (anonymous):

you were asked for the inverse right? not to actually divide 1 by 3 + i

OpenStudy (anonymous):

which is a totally different question